OpenStudy (anonymous):

Express the number in the term of i: 8√-4

OpenStudy (anonymous):

You can split this up as\[8\times \sqrt{4}\times \sqrt{-1}\]

OpenStudy (anonymous):

What is the square root of 4?

OpenStudy (jdoe0001):

hint: \(\bf \sqrt{-4}\implies \sqrt{4\cdot -1}\implies \sqrt{4}\cdot \sqrt{-1}\)

OpenStudy (anonymous):

2

OpenStudy (anonymous):

Then multiply that by 8...?

OpenStudy (anonymous):

16

OpenStudy (anonymous):

and what do you add at the end?

OpenStudy (anonymous):

so 16i?

OpenStudy (anonymous):

Yep :)

OpenStudy (anonymous):

Thanks! Also I have one quick question if you can answer it @Brill

OpenStudy (anonymous):

Sure go ahead

OpenStudy (anonymous):

One question was -√-32, I ended up with -2√8i. Am I correct..?

OpenStudy (anonymous):

Umm No

OpenStudy (anonymous):

Can you show me the steps?

OpenStudy (anonymous):

So split them up \[-1\times \sqrt{32}\times \sqrt{-1}\]

OpenStudy (anonymous):

What is the square root of 32?

OpenStudy (anonymous):

2√8

OpenStudy (anonymous):

Go on sqrt8 can be simplified more

OpenStudy (anonymous):

8√2?

OpenStudy (anonymous):

not quite Let's split up the expression\[2\sqrt{4}\sqrt{2}\]

OpenStudy (anonymous):

What is the square root of 4?

OpenStudy (anonymous):

2

OpenStudy (anonymous):

And 2 times 2?

OpenStudy (anonymous):

4

OpenStudy (anonymous):

So it's \[4\sqrt{2}\] right?

OpenStudy (anonymous):

And then what's square root of -1?

OpenStudy (anonymous):

Oh. I was adding instead of multiplying sorry :P

OpenStudy (anonymous):

i

OpenStudy (anonymous):

Oh yeah lol So what's your answer now?

OpenStudy (anonymous):

4√2i?

OpenStudy (anonymous):

Well yes but to avoid confusion the i is usually kept next to the 4

OpenStudy (anonymous):

Good job :)

OpenStudy (anonymous):

Thanks. Am I allowed to open up another random question to give you a medal? I mean you earned it

OpenStudy (anonymous):

Nah it's fine I don't really need it :)

OpenStudy (anonymous):

Haha thanks :)